OpenStudy (anonymous):

Given cos 0=-4/5, and the angle 0 is in quadrant two, find the value for sin 0.

OpenStudy (blockcolder):

You can use Pythagorean identity w/ sine and cosine to find \(\sin\theta\) given \(\cos\theta\) then use the fact that \(\theta\) is in Quadrant 2.

OpenStudy (anonymous):



OpenStudy (anonymous):

the 3 via pythagoras or more likely just remembering the famous 3 - 4 - 5 right triangle then sine is opposite over hypotenuse, and it is positive since you are in quadrant 2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw